OpenStudy (anonymous):

looking to find the differential (dy) when (x=1) and (dx=0.1) for the equation y=4x^2=3x=3 Step by step would be great?

OpenStudy (anonymous):

\[y=4x^2+3x+3\]?

OpenStudy (amistre64):

\[\int dy = (4\int (x^2)dx)+(3\int(x)dx)+(3\int dx)\]

OpenStudy (anonymous):

\[y = 4x ^{2}+3x+3\] to find y' the derivative can be spread among the different additive portions then using the Power rule on each variable you find \[y' = (4x^2)'+(3x)'+3\] \[y' = (4*2)x+(3*1)+0\] where the derivative of a constant = 0

OpenStudy (anonymous):

confused by the equal signs. guru?

OpenStudy (amistre64):

nah master, its a common typo on my end too lol

OpenStudy (anonymous):

i think that is wrong. you are just looking for the differential yes?

OpenStudy (anonymous):

amistre64 the question is only asking for the derivative, not the integral ;-)

OpenStudy (amistre64):

lol ..... well, when the question to my answer comes around; just point them to it ;)

OpenStudy (amistre64):

diffrential; differential equations; ... they really should come up with more words to use for this stuff ;)